Light enters train air to glass having refractive index 1.50. What is the speed of light in the glass? The speed ref light in vacuum is 3 × 108 m/s.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


The speed of light in vacuum is = 3 × 108m/s
Refractive index of glass n = 1.50


n = c / v
Speed of light in Glass, v = c / n
v = 2 * 10^8 m/s

Thanks